OpenStudy (ineedhelp10):

A square and a semicircular region have the same perimeter. If the length of the radius of the semicircular region is 16, what is the length of one side of the square?

OpenStudy (ineedhelp10):

@Hero

OpenStudy (ineedhelp10):

@Luigi0210

hartnn (hartnn):

perimeter (circumference) of circle is ?

hartnn (hartnn):

you know ?

OpenStudy (anonymous):

A semi circle is half circle

OpenStudy (ineedhelp10):

pi(radius)^2

OpenStudy (anonymous):

That is the area

hartnn (hartnn):

thats the area of circle i asked for circumference

OpenStudy (anonymous):

Perimeter = 2*pi*radius

OpenStudy (ineedhelp10):

circumference(diameter)

hartnn (hartnn):

you know what circumference is ?

OpenStudy (anonymous):

Do u mean pi*diameter

OpenStudy (ineedhelp10):

oh yeah @UditKulka

OpenStudy (anonymous):

well then that's full circle what is for half circle (semi-circle)?

OpenStudy (anonymous):

well then that's full circle what is for half circle (semi-circle)? @ineedhelp10

OpenStudy (ineedhelp10):

wouldny i just need to divide it by to get the semicircle?

hartnn (hartnn):

by 2 , yes

OpenStudy (anonymous):

Divide it by 2 + the diameter

OpenStudy (anonymous):

So, pi*r + 2r r= radius 2r gives the diameter right????

OpenStudy (anonymous):

perimeter of semicircular region = \[\pi r + 2r = 3.14 \times 16 + \times 16 =50.24+32= 82.24 \] But according to the question perimeter of square = perimeter of semicircular region i.e. perimeter of square = 82.24 units let the side of the square be a 4a = 82.24\[a= \frac{82.24}{4}= 20.56\] Hence length of the side of the square is 20.56 units @ineedhelp10

OpenStudy (anonymous):

@hartnn sory for giving full solution..

OpenStudy (ineedhelp10):

so the perimeter formula for a semicircle is pi(r)+2r? @dpasingh

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ineedhelp10 yes dear